# Depth-Breadth Synergy in RLVR: Unlocking LLM Reasoning Gains with Adaptive Exploration

### v5

@@ 1 Introduction
However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K. In this paper, we conduct a systematic analysis of these two under-exploited dimensions in RLVR: Depth and Breadth.
@@
  • We introduce DARS, which reallocates compute from medium difficulty problems to the hardest problems via multi-stage rollout sampling. DARS re-weights the cumulative advantage distribution and quantitatively expands the sparse reward signals for difficult problems. In practice, DARS significantly improves Pass@K over multiple benchmarks.
  • We further illustrate that large breadth in RLVR training matters for the Pass@1 performance. Moreover, by combining DARS with large breadth training, we reveal the complementarity of Depth and Breadth in RLVR and acquire simultaneous boosts in both Pass@K and Pass@1 performance.
@@ Appendix A More Related Works
Appendix B Derivation of Additional Rollouts $\Delta n_{j}$
@@
Appendix C Mathematical Derivation: Group Cumulative Advantage and GRPO Gradient Norm
In this appendix, we provide a detailed mathematical derivation demonstrating the relationship between Group Cumulative Advantage and the gradient norm in GRPO. This derivation substantiates the claim that $\mathcal{A}_{\text{group}}$ serves as an effective indicator of the model’s attention to specific problems during training.
C.1 GRPO Gradient Formulation
The GRPO objective function and its gradient are given by (we can remove the clip operation for simplified analysis):
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\bigg],$ (11)
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\nabla_{\theta}\rho_{i,t}(\theta)\bigg]$
$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\frac{\nabla_{\theta}\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}\bigg]$
$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\bigg].$ (12)
C.2 Derivation of Gradient Norm Upper Bound
We now derive the upper bound relationship between the gradient norm and Group Cumulative Advantage.
The gradient norm of our adopted GRPO algorithm is shown as the following:
@@ C.2 Derivation of Gradient Norm Upper Bound
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|=\left\|\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right]\right\|.$ (13)
Since the norm is a convex function and expectation is linear, by Jensen’s inequality:
$\left\|\mathbb{E}[X]\right\|\leq\mathbb{E}[\|X\|]$ (14)
Thus:
@@
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\left\|\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (15)
Applying the triangle inequality to the inner summation:
$\left\|\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}a_{i,t}\right\|\leq\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|a_{i,t}\right\|,$ (16)
where:
$a_{i,t}=\frac{1}{\sum_{i=1}^{G}|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t}).$ (17)
Therefore:
@@
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (18)
Then we have,
$\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|=|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|.$ (19)
Thus:
@@
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (20)
We further take a boundedness assumption. in policy optimization, we assume the gradient log-probabilities $\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|$ are bounded. Furthermore, it’s vital to notice that the importance ratio $\rho_{i,t}(\theta)$ are also bounded through the clip operation in GRPO algorithm.
Thus, there exists a constant $C>0$ such that:
$|\rho_{i,t}(\theta)|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\leq C.$ (21)
Therefore:
@@
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|\right].$ (22)
Noting that $|\hat{A}_{i}|$ is independent of $t$ for fixed $i$:
$\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|=\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|.$ (23)
Thus:
@@
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\right].$ (24)
In this paper, we define the Group Cumulative Advantage as:
$\mathcal{A}_{\text{group}}=\sum_{i=1}^{G}|\hat{A}_{i}|.$ (25)
By the weighted arithmetic mean inequality:
$\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\leq\frac{\sum_{i=1}^{G}|o_{i}|\cdot\sum_{i=1}^{G}|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}=\sum_{i=1}^{G}|\hat{A}_{i}|=\mathcal{A}_{\text{group}}.$ (26)
Therefore:
$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\mathcal{A}_{\text{group}}\right].$ (27)
C.3 Theoretical Implications
This derivation establishes that Group Cumulative Advantage $\mathcal{A}_{\text{group}}$ provides an upper bound for the expected gradient norm in GRPO. Consequently: Larger $\mathcal{A}_{\text{group}}$ values indicate stronger gradient signals for the corresponding problem. The training process allocates more ”attention” to problems with higher $\mathcal{A}_{\text{group}}$ values during parameter updates. Therefore, we consider $\mathcal{A}_{\text{group}}$ serves as a mathematically grounded indicator of problem importance in GRPO training dynamics.
This theoretical foundation validates the use of Group Cumulative Advantage as a meaningful metric for analyzing training behavior and problem prioritization in GRPO.

### v6

@@ 1 Introduction
However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K. In this paper, we conduct a systematic analysis of these two under-exploited dimensions in RLVR.
@@
  • We introduce DARS, which reallocates compute to the hard problems via multi-stage rollout sampling. DARS re-weights the cumulative advantage distribution and quantitatively expands the sparse reward signals for difficult problems. In practice, DARS significantly improves Pass@K performance. We elaborate on the impact of cumulative advantage on the gradient norm in Appendix B.2, and further discuss the relationship between DARS-HW and MaxRL (Tajwar et al., 2026) in Appendix C.
  • We further illustrate that large breadth matters for the Pass@1 performance. Moreover, by combining DARS with large breadth training, we reveal the complementarity of Depth and Breadth in RLVR and acquire simultaneous boosts in both Pass@K and Pass@1 performance.
@@ Appendix A More Related Works
Appendix B Mathematical Derivations for DARS and GRPO Training Dynamics
B.1 Derivation of Additional Rollouts $\Delta n_{j}$
@@
B.2 Connection of Expected Squared Gradient Norm and Group Cumulative Advantage
In this subsection, we will prove that: in the setting of group-relative advantages, the expected squared gradient norm allocated to a problem is proportional to the group cumulative advantage $\mathcal{A}_{\text{group}}$.
Proposition B.1 In a Reinforcement Learning with Verifiable Reward (RLVR) setting utilizing group-relative advantages, the expected squared gradient norm of a problem group is directly proportional to its Group Cumulative Advantage $\mathcal{A}_{\text{group}}$.
Proof.
By definition, for a specific problem $q$, the group gradient $\mathbf{g}_{\text{group}}$ before taking the dataset expectation is formulated at the trajectory level as:
$\mathbf{g}_{\text{group}}=\frac{1}{G}\sum_{i=1}^{G}\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i}|q),$ (11)
where $o_{i}$ denotes the $i$-th sampled trajectory. Let $\mathbf{g}_{i}\triangleq\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)$ denote the score function of the trajectory. To evaluate the optimization step size, we compute the expected squared $L_{2}$ norm of the group gradient:
$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]=\frac{1}{G^{2}}\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\|\mathbf{g}_{i}\|^{2}+\sum_{i\neq j}\hat{A}_{i}\hat{A}_{j}\langle\mathbf{g}_{i},\mathbf{g}_{j}\rangle\right].$ (12)
While the cross terms $\mathbb{E}[\hat{A}_{i}\hat{A}_{j}\langle\mathbf{g}_{i},\mathbf{g}_{j}\rangle]$ do not strictly vanish (since the advantages $\hat{A}_{i}$ and $\hat{A}_{j}$ are coupled through the group mean reward and correlated with the score functions), they are heavily dominated by the diagonal terms in the context of Large Language Models (LLMs).
In the enormously high-dimensional parameter space typical of LLMs, the policy gradient estimator exhibits notoriously high variance. The individual trajectory gradients $\mathbf{g}_{i}$ are highly stochastic and point in nearly orthogonal directions. Consequently, the inner product of different trajectory gradients $\langle\mathbf{g}_{i},\mathbf{g}_{j}\rangle$ is exceedingly small compared to the squared norm of a single trajectory gradient $\|\mathbf{g}_{i}\|^{2}$. Mathematically, the diagonal sum represents the uncentered variance (update energy) of the samples, while the cross-terms correspond to the squared norm of the expected gradient $\|\mathbb{E}[\mathbf{g}_{\text{group}}]\|^{2}$. In RL for LLMs, the gradient noise variance inherently dwarfs the true gradient magnitude (i.e., $\text{Tr}(\text{Var}(\hat{A}_{i}\mathbf{g}_{i}))\gg\|\mathbb{E}[\hat{A}_{i}\mathbf{g}_{i}]\|^{2}$). Therefore, we can safely approximate the expected squared group gradient by its dominant diagonal components:
@@ B.2 Connection of Expected Squared Gradient Norm and Group Cumulative Advantage
$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]\approx\frac{1}{G^{2}}\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\|\mathbf{g}_{i}\|^{2}\right].$ (13)
Let $V^{2}=\mathbb{E}[\|\mathbf{g}_{i}\|^{2}]$ denote the expected squared norm of the trajectory gradient (which corresponds to the trace of the Fisher Information Matrix). Under the standard simplifying assumption that the gradient magnitude is approximately independent of the specific binary correctness $r_{i}$, we can decouple the expectations. This elegantly isolates the advantage term, simplifying the expected squared group gradient to:
@@
$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]\approx\frac{V^{2}}{G^{2}}\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\right].$ (14)
We now evaluate the advantage term $\sum_{i=1}^{G}\hat{A}_{i}^{2}$. In RLVR, verifiable rewards are binary, i.e., $r_{i}\in\{0,1\}$. Without standard-deviation normalization, the advantage is computed as $\hat{A}_{i}=r_{i}-u$, where $u$ is the expected group mean reward. The sum of squared advantages expands algebraically as:
$\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\right]$ $=\sum_{i:r_{i}=1}(1-u)^{2}+\sum_{i:r_{i}=0}(-u)^{2}$
$=c(1-u)^{2}+(G-c)u^{2}$
$=Gu(1-u)^{2}+G(1-u)u^{2}$
$=Gu(1-u)[(1-u)+u]=Gu(1-u).$ (15)
Simultaneously, the Group Cumulative Advantage $\mathcal{A}_{\text{group}}$ is defined as the sum of absolute advantages:
$\mathbb{E}\left[\mathcal{A}_{\text{group}}\right]$ $=\sum_{i=1}^{G}|\hat{A}_{i}|=\sum_{i:r_{i}=1}|1-u|+\sum_{i:r_{i}=0}|-u|$
$=c(1-u)+(G-c)u$
$=Gu(1-u)+G(1-u)u=2Gu(1-u).$ (16)
Comparing Equation 15 and Equation 16 reveals a strict algebraic identity unique to the binary group-relative advantage formulation:
@@
$\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\right]\equiv\frac{1}{2}\mathbb{E}\left[\mathcal{A}_{\text{group}}\right].$ (17)
Substituting Equation 17 into Equation 14, we obtain the final exact proportionality:
@@
$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]\approx\left(\frac{V^{2}}{2G^{2}}\right)\mathbb{E}[\mathcal{A}_{\text{group}}]=2Gu(1-u).$ (18)
This demonstrates that the expected optimization step size (squared gradient norm) is strictly proportional to the Group Cumulative Advantage. ∎
Theoretical Implications. Proposition B.2 fundamentally justifies our algorithmic design. By demonstrating that the expected variance (i.e., update energy) of the parameter gradient is mathematically locked to $\mathcal{A}_{\text{group}}$, we confirm that $\mathcal{A}_{\text{group}}$ acts as a rigorous surrogate for the gradient signal strength allocated to each problem. This mathematically verifies that existing GRPO variants intrinsically under-update hard problems (where $\mathcal{A}_{\text{group}}$ is distortedly low), directly motivating the necessity of our Difficulty Adaptive Rollout Sampling (DARS) methodology to re-balance update magnitudes.
Appendix C Theoretical Connection Between DARS and Maximum Likelihood Reinforcement Learning
Recent work on Maximum Likelihood Reinforcement Learning (MaxRL) (Tajwar et al., 2026) demonstrates that standard expected-reward objectives systematically under-weight high-difficulty instances. To counteract this, MaxRL introduces a modified advantage function to approximate the Maximum Likelihood (ML) objective.
In this section, we provide a formal theoretical analysis demonstrating that our Difficulty-Adaptive Rollout Sampling (DARS) with the Hardness-Weighted (HW) schedule optimizes an expected objective identical to MaxRL. Crucially, we prove that while MaxRL relies on algebraic advantage scaling, which inherently exacerbates gradient variance. While DARS achieves the same objective through physical compute reallocation. We establish that DARS serves as a strict variance-reduced estimator of the ML-equivalent objective, offering superior optimization stability.
C.1 Objective Equivalence via Cumulative Advantage
We first analyze the weighting behavior of both algorithms utilizing our proposed Group Cumulative Advantage ($\mathcal{A}_{group}$) framework. Let $N$ denote the number of rollouts for a given prompt $q$, and let $u\in(0,1)$ denote the mean reward (i.e., empirical accuracy or pass rate), where the binary reward is $r_{i}\in\{0,1\}$.
Proposition C.1 (Equivalence of Difficulty Weighting Profile) Both MaxRL and DARS-HW impose a problem-level optimization weight that is strictly proportional to (1u)(1-u). Consequently, both algorithms share an identical inductive bias: linearly up-weighting the optimization focus on problems as their difficulty increases ($u\to 0$).
Proof.
For MaxRL, the advantage is scaled as $\hat{A}^{MaxRL}_{i}=\frac{r_{i}-u}{u}$. For a group of $N$ rollouts, the expected number of successful trajectories ($r_{i}=1$) is $Nu$, and the expected number of failed trajectories ($r_{i}=0$) is N(1u)N(1-u). The cumulative advantage is thus computed as:
$\mathcal{A}_{group}^{MaxRL}(u)$ $=\mathbb{E}\left[\sum_{i=1}^{N}\left|\hat{A}^{MaxRL}_{i}\right|\right]$
$=Nu\cdot\left|\frac{1-u}{u}\right|+N(1-u)\cdot\left|\frac{0-u}{u}\right|$
$=N(1-u)+N(1-u)=2N(1-u)$
For DARS-HW, we employ the unscaled advantage $\hat{A}_{i}=r_{i}-u$. Based on our proposed schedule (Equation 7), the target cumulative advantage is explicitly constrained to:
$\mathcal{A}_{group}^{HW}(u)=2(1-u)\mathcal{A}_{group}^{N_{pre}}(0.5)=C\cdot(1-u)$
where $C=2\mathcal{A}_{group}^{N_{pre}}(0.5)$ is a constant independent of the prompt difficulty $u$. Since $\mathcal{A}_{group}^{MaxRL}(u)\propto(1-u)$ and $\mathcal{A}_{group}^{HW}(u)\propto(1-u)$, the weighting profiles are equivalent. ∎
C.2 Equivalence of Expected Gradients
To formalize the alignment in optimization trajectories, we compare the total expected gradient contribution of a single prompt $q$ under both methods. Let $\nabla J_{i}(u)=(r_{i}-u)\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)$ denote the standard baseline-subtracted policy gradient for a single rollout.
Proposition C.2 (Equivalence of Expected Gradient Updates) Let $K$ be a positive scalar. The expected gradient updates of MaxRL and DARS-HW are proportional: $\mathbb{E}[\nabla J^{MaxRL}_{q}]\propto\mathbb{E}[\nabla J^{DARS-HW}_{q}]\propto\frac{1}{u}\mathbb{E}[\nabla J_{i}(u)]$. Both methods approximate the Maximum Likelihood gradient by universally scaling the standard policy gradient by a factor of $1/u$.
Proof.
For MaxRL, using a fixed rollout size $N$, the total expected gradient is scaled algebraically:
@@ C.2 Equivalence of Expected Gradients
$\mathbb{E}[\nabla J^{MaxRL}_{q}]=\mathbb{E}\left[\sum_{i=1}^{N}\frac{r_{i}-u}{u}\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\right]=\frac{N}{u}\cdot\mathbb{E}\left[\nabla J_{i}(u)\right]$
For DARS-HW, we dynamically adjust the rollout size $N_{DARS}(u)$ to match the target $\mathcal{A}_{group}^{HW}(u)=C(1-u)$. Recall the standard group cumulative advantage property $\mathcal{A}_{group}=2N_{DARS}u(1-u)$. Equating the two yields $2N_{DARS}(u)u(1-u)=C(1-u)$, which simplifies to $N_{DARS}(u)=\frac{C}{2u}$. Thus, the expected gradient contribution is:
@@
$\mathbb{E}[\nabla J^{DARS-HW}_{q}]=\mathbb{E}\left[\sum_{i=1}^{N_{DARS}(u)}(r_{i}-u)\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\right]=\frac{C}{2u}\cdot\mathbb{E}\left[\nabla J_{i}(u)\right]$
Up to the constants $N$ and $C/2$, both expected gradients scale strictly by $1/u$. ∎
C.3 Variance Analysis: DARS-HW versus MaxRL
While Proposition C.2 establishes that DARS-HW and MaxRL optimize the same expected objective, they diverge significantly in their construction of the Monte Carlo gradient estimator. Let $\Sigma_{g}=\text{Var}(\nabla J_{i}(u))$ denote the trace of the covariance matrix of the unscaled token-level gradient for a single rollout. We make the standard assumption that conditioned on the prompt $q$ and current policy $\pi_{\theta}$, individual rollouts are independent and identically distributed (i.i.d.).
Let $u\to 0$ denote the regime of increasingly difficult reasoning tasks. For the MaxRL estimator, the variance relies on algebraic scalar multiplication over a fixed $N$ samples. By the properties of variance for i.i.d. variables ($\text{Var}(aX)=a^{2}\text{Var}(X)$):
$\text{Var}(\hat{\nabla}J^{MaxRL}_{q})=\text{Var}\left(\sum_{i=1}^{N}\frac{1}{u}\nabla J_{i}(u)\right)=\sum_{i=1}^{N}\frac{1}{u^{2}}\text{Var}(\nabla J_{i}(u))=\frac{N}{u^{2}}\Sigma_{g}$
Conversely, DARS-HW physically expands the sampling space, drawing $N_{DARS}(u)=\frac{C}{2u}$ independent trajectories while maintaining the bounded unscaled advantage $|r_{i}-u|\leq 1$:
$\text{Var}(\hat{\nabla}J^{DARS-HW}_{q})=\text{Var}\left(\sum_{i=1}^{N_{DARS}(u)}\nabla J_{i}(u)\right)=\sum_{i=1}^{N_{DARS}(u)}\text{Var}(\nabla J_{i}(u))=N_{DARS}(u)\Sigma_{g}=\frac{C}{2u}\Sigma_{g}$
Taking the ratio of the two variances yields:
$\frac{\text{Var}(\hat{\nabla}J^{MaxRL}_{q})}{\text{Var}(\hat{\nabla}J^{DARS-HW}_{q})}=\frac{N/u^{2}}{C/(2u)}=\frac{2N}{C}\cdot\frac{1}{u}\propto\frac{1}{u}$
Therefore, as $u\to 0$, the variance of MaxRL grows at a rate of $\mathcal{O}(1/u^{2})$, while DARS-HW grows at a significantly slower rate of $\mathcal{O}(1/u)$.
C.4 Theoretical Analysis Conclusions
  • Unified Optimization Objective with Divergent Implementation Routes. DARS-HW and MaxRL are two distinct implementation routes toward the same Maximum Likelihood reinforcement learning objective. MaxRL achieves the target optimization weight by algebraically modifying the advantage function ($\hat{A}^{MaxRL}_{i}=\frac{r_{i}-u}{u}$), scaling the gradient signal of each rollout by $1/u$. In contrast, DARS-HW retains the original unscaled advantage function ($\hat{A}_{i}=r_{i}-u$) and instead achieves the identical expected gradient scaling by adaptively increasing the number of rollouts ($N_{DARS}(u)\propto 1/u$) for difficult problems. These two approaches are mathematically equivalent in expectation but differ fundamentally in their implementation strategy.
  • Variance Reduction of DARS. Appendix C.3 fundamentally explains the superiority of DARS over pure advantage-scaling methods as MaxRL. For complex reasoning tasks where the initial pass rate is extremely low ($u\ll 1$), MaxRL heavily amplifies the gradient of rare, potentially idiosyncratic correct trajectories. This leads to a high-variance gradient regime, which may inducing optimization instability and premature convergence. By contrast, DARS-HW functions as a variance-reduced estimator. Instead of amplifying a small fixed set of rollouts algebraically, DARS-HW allocates explicit computational budget to difficult problems, thereby expanding the breadth of actual correct reasoning paths discovered. This mechanism preserves a high signal-to-noise ratio in the gradient and maintains robust token-level entropy, theoretically validating DARS-HW as a more stable and efficient paradigm for Maximum Likelihood RL optimization.